OpenStudy (anonymous):

Medal Given!!! If y − 3 = 3x, which of the following sets represents possible inputs and outputs of the function, represented as ordered pairs?

OpenStudy (danjs):

y = 3*x + 3

OpenStudy (danjs):

It is a LINE

OpenStudy (anonymous):

(0.3) (1.6) (2.9) (3.0) (6.1) (9.2) (1.3) (2.6) (3.9) (3.1) (6.2) (9.3)

OpenStudy (danjs):

Each of those points is (x,y) use the x and see if y=3x+3

OpenStudy (anonymous):

b?

OpenStudy (elonasushchik):

Correct!

OpenStudy (danjs):

(x,y) = (0,3) 3 = 3(0) + 3 true

OpenStudy (anonymous):

thank you @ElonaSushchik and @DanJS .

OpenStudy (danjs):

no B is not correct, the first point doesnt work (x,y) = (3,0) 0 = 9 False

OpenStudy (elonasushchik):

Medal?!

OpenStudy (danjs):

are you even on the same question elona, what is this about a parabola?

OpenStudy (anonymous):

c?

OpenStudy (elonasushchik):

Oops wrong person. Lol sorry!

OpenStudy (danjs):

A works, you have to test all 3 points in the equation to see if it is true

OpenStudy (anonymous):

@DanJS thank you so much for helping me!

OpenStudy (danjs):

Just dont guess the answer. In C the first point is (x,y) = (1,3) y-3 = 3x ? 3 - 3 = 3*1 0 = 3 FALSE

OpenStudy (anonymous):

@DanJS Thank you! i understand the question now.

OpenStudy (danjs):

k, np

OpenStudy (anonymous):

:) Merry Christmas!